OpenStudy (anonymous):

5(x - 3) + 4(x + 3) = 3(x - 1) solve for x

OpenStudy (anonymous):

first reorder the terms: 5(-3 + x) + 4(x + 3) = 3(x + -1) (-3 * 5 + x * 5) + 4(x + 3) = 3(x + -1) (-15 + 5x) + 4(x + 3) = 3(x + -1)